OpenStudy (anonymous):

96 miles in 1.5 hours.What is the speed in miles per hour? A.48 miles per hour B.54 miles per hour C.64 miles per hour D.144 miles peer hour

OpenStudy (cwrw238):

96 / 1.5

OpenStudy (anonymous):

the answer is C

OpenStudy (anonymous):

the formula for speed is speed = distance / time so speed = 96/1.5 solve and u gettit !!

OpenStudy (whpalmer4):

this is a rate problem: rate = \(\large\frac{96\text{ miles}}{1.5 \text{ hours}} = \frac{96}{1.5}\frac{\text{ miles}}{\text{ hour}} = \text{______} \frac{\text{miles}}{\text{hour}}\)